# Useful stochastic bounds in time-varying queues with service and patience times having general joint distribution

### v1

@@ Abstract
where $W_{x}(\cdot)$ is the workload process in the queue and $g(\cdot)$ is any lower semi-continuous function. We also demonstrate how to utilise these bounds via some examples under the additional assumption that $\lambda(\cdot)$ is periodic.
@@ 1 Introduction
$G(\cdot)\equiv G(\Psi)(\cdot)\equiv\Psi(\cdot,\infty)\equiv\lim_{y\uparrow\infty}\Psi(\cdot,y)\,,$ (1)
@@
$\ H(\cdot)\equiv H(\Psi)(\cdot)\equiv\Psi(\infty,\cdot)\equiv\lim_{s\uparrow\infty}\Psi(s,\cdot)\,.$ (2)
In particular, when $\Psi(\cdot)$ has a product-form (i.e., $\Psi(s,y)\equiv G(s)H(y)$ for any $(s,y)\in\mathbb{R}^{2}$), this model is known in literature as the $\text{M}_{t}/\text{G}/1+\text{H}$ queue (see, the work by Bodas, Mandjes and Ravner [10] for a multi-server version of this model). If, in addition, $\lambda(\cdot)$ is a constant function, then this queue will become M/G/$1$+H queue which have already been studied by several authors, e.g., Baccelli, Boyer and Hebuterne [4], Bae and Kim [5], Perry and Asmussen [47], Boxma, Perry, and Stadje [12] and Boxma, Perry, Stadje and Zacks, [11].
@@ 1.1 General description of the problem
However, as illustrated by the lower drawing of Figure 1, if the customers are allowed to have finite patience times, then (3) becomes false. This observation leads to the following question: Is there some sense (weaker than the one in (3)) in which $W_{\lambda}$ is stochastically dominated by $W_{\lambda_{h}}$? In the present work we provide an affirmative answer for this query by specifying a general a class of functionals $\mathcal{H}$ such that
@@
where $\leq_{\text{st}}$ stands for a first-order stochastic dominance between the distributions of two random variables (see, e.g., the book by Shaked and Shanthikumar [43]).
@@ 1.2 The main results
$A_{\lambda}(g)\equiv A_{\lambda}(g;x)\equiv A_{\lambda}(g;x,\Psi)\equiv\int_{0}^{\tau_{\lambda}(x;\Psi)}g\circ W_{\lambda}(t;x,\Psi){\rm d}t\,,$ (6)
@@
i.e., the set of all non-negative lower semi-continuous functions on $[0,\infty)$. The next Theorems 1 and 2 below are the main results of this work as they jointly construct a general class of functionals for which (4) is valid.
@@
$A_{\lambda}(g;x,\Psi)\leq_{\text{st}}A_{\lambda_{h}}(g;x,\Psi),\ \forall x>0,\ g\in\mathcal{G},\ \Psi\in\mathcal{L}.$ (8)
@@
$A^{*}_{\lambda}(g;x,\Psi)\leq_{\text{st}}A^{*}_{\lambda_{h}}(g;x,\Psi),\ \forall x>0,\ g\in\mathcal{G},\ \Psi\in\mathcal{L}\,.$ (10)
@@ 1.3 Periodic $\text{M}_{t}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queues and applications
$\xi_{\lambda}\equiv\inf\left\{\kappa n;\ W_{\lambda}(\kappa n)=0,\ n\in\mathbb{N}\right\}\,.$ (12)
@@
$\mathcal{A}^{*}_{\lambda}(g)\equiv\mathcal{A}^{*}_{\lambda}(g;\Psi)\equiv\int_{0}^{\xi_{\lambda}}g\circ W_{\lambda}(t-){\rm d}N_{\lambda}^{*}(t)\,.$ (14)
@@
Assume that $g\in\mathcal{G}$, $\Psi\in\mathcal{L}$ and $\lambda(\cdot)$ has a period $\kappa>0$. In addition, let
$F_{\lambda}(\cdot)\equiv F_{\lambda}(\cdot;g,\Psi)\equiv{\text{dist}}\left[A_{\lambda}(g;x,\Psi)\right],$ (15)
$F^{*}_{\lambda}(\cdot)\equiv F^{*}_{\lambda}(\cdot;g,\Psi)\equiv{\text{dist}}\left[A_{\lambda}^{*}(g;x,\Psi)\right],$
when the initial distribution of $x$ is $G(\cdot)\equiv G(\Psi)(\cdot)$. In addition, denote the next two geometric decompoundings of $F_{\lambda}(\cdot)$ and $F^{*}_{\lambda}(\cdot)$:
$J_{\lambda_{h}}(u;g,\Psi)\equiv\sum_{n=1}^{\infty}\left(1-e^{-\kappa\lambda_{h}}\right)^{n-1}e^{-\kappa\lambda_{h}}F^{\star n}_{\lambda}(u-\max\{1,\kappa\}),\ \forall u\in\mathbb{R}\,,$ (16)
$J^{*}_{\lambda_{h}}(u;g,\Psi)\equiv\sum_{n=1}^{\infty}\left(1-e^{-\kappa\lambda_{h}}\right)^{n-1}e^{-\kappa\lambda_{h}}\left(F^{*}_{\lambda}\right)^{\star n}(u-\max\{1,\kappa\}),\ \forall u\in\mathbb{R}\,.$
@@
${\text{dist}}\left[\mathcal{A}_{\lambda}(g;\Psi)\right]\leq_{\text{st}}J_{\lambda_{h}}(\cdot;g,\Psi)\ \ \text{and}\ \ {\text{dist}}\left[\mathcal{A}^{*}_{\lambda}(g;\Psi)\right]\leq_{\text{st}}J^{*}_{\lambda_{h}}(\cdot;g,\Psi)\,.$ (17)
@@
$\lim_{u\to\infty}\sup\frac{\mathbb{P}\left\{\xi_{\lambda}>u\right\}}{1-G\left\{[u-g(0)\max\{\kappa,1\}](1-\rho_{h})\right\}}\leq\frac{1-e^{-\kappa\lambda_{h}}}{(1-\rho_{h})e^{-\kappa\lambda_{h}}}\,.$ (19)
@@
De Meyer and Teugels [14] proved that (18) is valid whenever $G(\cdot)$ is regularly-varying. Later, Zwart [55] extended this result to the class of all intermediate regularly-varying. It turns out that (18) is valid also when the tail of $G(\cdot)$ is heavier than $e^{-\sqrt{u}}$ (for the precise details see the works by Jelenković, P. R., and Momčilović [28] and Baltrūnas, Daley and Klüppelberg [6]). Another related work is by Asmussen, Klüppelberg and Sigman [3], who explained why (18) cannot hold when the tail of $G(\cdot)$ is lighter than $e^{-\sqrt{u}}$.
@@
$\lim_{u\to\infty}\sup\frac{\mathbb{P}\left\{\eta^{*}_{\lambda}>u\right\}}{1-G\left\{\frac{\left[u-g(0)\max\{\kappa,1\}\right](1-\rho_{h})}{\lambda_{h}}\right\}}\leq\frac{1-e^{-\kappa\lambda_{h}}}{(1-\rho_{h})e^{-\kappa\lambda_{h}}}\,.$ (21)
@@
Note that the fact that the $\text{M}_{t}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queue is stable under the assumption that $\Psi(\cdot)$ has a product-form and (22) is valid can also be proven by applying Lyapunov functions, see, e.g., Theorem 4.1 (and its proof) in the recent work by Bodas, Mandjes and Ravner [10].
@@
$\mathbb{E}W_{\lambda}(\infty)=\frac{\mathbb{E}\mathcal{A}_{\lambda}(g)}{\mathbb{E}\xi_{\lambda}}\,,$ (23)
@@ 2 Proofs of Theorems 1 and 2
In this section we supply the proofs of the main results of the present work (Theorem 1 and Theorem 2). Since the proofs of these two theorems are essentially identical (up to a few minor modifications), we will only provide the detailed proof of Theorem 1.
2.1 General guidelines of the proof
The first step in the proof of Theorem 1 is to define the following auxiliary queue:
Definition 2
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    1. (a)
      • There is empty place for him, i.e., the left limit of the queue-length process (counting also the potential customer in the service position) at his arrival epoch is less than $k+1$.
    2. (b)
Remark 4
$\tau^{(k)}_{\lambda}(x)\equiv\tau^{(k)}_{\lambda}\left(x;\Psi\right)\equiv\inf\left\{t\geq 0;W^{(k)}_{\lambda}\left(t;x,\Psi\right)=0\right\},$ (24)
and also for any Borel function $g:[0,\infty)\rightarrow[0,\infty)$
$A^{(k)}_{\lambda}(g)\equiv A^{(k)}_{\lambda}(g;x)\equiv A^{(k)}_{\lambda}(g;x,\Psi)\equiv\int_{0}^{\tau^{(k)}_{\lambda}(x;\Psi)}g\circ W^{(k)}_{\lambda}(t;x,\Psi){\rm d}t\,,$ (25)
Below, we will prove:
Proposition 1
Assume that $\lambda(t)\leq\lambda_{h}$ for every $t\geq 0$. Then, for every Borel function $g:[0,\infty)\rightarrow[0,\infty)$, $x>0$, $\Psi\in\mathcal{L}$ and $k\in\mathbb{Z}_{\geq 0}$, the stochastic dominance $A^{(k)}_{\lambda}\left(g;x,\Psi\right)\leq_{\text{st}}A^{(k)}_{\lambda_{h}}\left(g;x,\Psi\right)$ holds.
@@ 2.2 Proof of Proposition 1
As announced, we use an inductive argument.
@@
$W_{\lambda}^{(0)}(t;x,\Psi)=W_{\lambda_{h}}^{(0)}(t;x,\Psi)=x-t\ \ ,\ \ \forall\ t\geq 0\,,x>0\ ,\ \Psi\in\mathcal{L}\,,$ (26)
@@
$\int_{I_{i}}g\circ W_{\lambda}^{(k+1)}(t;x,\Psi){\rm d}t\stackrel{{\scriptstyle{\text{d}}}}{{=}}A^{(k)}_{\lambda(T_{i}+\cdot)}\left[g(\cdot+S_{i});D_{i},\Psi\left(\cdot,\cdot+S_{i}\right)\right]$ (28)
for $1\leq i\leq\eta$. Now, for each $1\leq i\leq\eta$ fix values $(d,s,t)\in\mathcal{S}\equiv\text{Supp}\left(D_{i},S_{i},T_{i}\right)$ and let $F(\cdot;d,s,t)$ (resp., $G(\cdot;d,s)$) be the cdf of
$A^{(k)}_{\lambda(t+\cdot)}\left[g(\cdot+s);d,\Psi\left(\cdot,\cdot+s\right)\right]\ \ \left(\text{resp., }A^{(k)}_{\lambda_{h}}\left[g(\cdot+s);d,\Psi\left(\cdot,\cdot+s\right)\right]\right).$ (29)
@@
$V_{i}(d,s,t)\equiv F^{-1}\left(U_{i};d,s,t\right),\qquad Z_{i}(d,s)\equiv G^{-1}\left(U_{i};d,s\right),$ (30)
@@
$V_{i}\equiv\begin{cases}V_{i}\left(D_{i},S_{i},T_{i}\right),&1\leq i\leq\eta,\\ 0,&\text{otherwise},\end{cases}\qquad\ Z_{i}\equiv\begin{cases}Z_{i}\left(D_{i},S_{i}\right),&1\leq i\leq\eta,\\ 0&\text{otherwise},\end{cases}$ (31)
@@
$\sum_{i=1}^{\eta}V_{i}\stackrel{{\scriptstyle\text{d}}}{{=}}\sum_{i=1}^{\eta}\int_{I_{i}}g\circ W_{\lambda}^{(k+1)}(t;x,\Psi){\rm d}t$ (33)
@@
$\sum_{i=1}^{\eta}\int_{I_{i}}g\circ W_{\lambda}^{(k+1)}(t;x,\Psi){\rm d}t\leq_{\text{st}}\sum_{i=1}^{\eta}Z_{i}\,,$ (34)
@@
$\int_{I_{i}(\lambda_{h})}g\circ W_{\lambda_{h}}^{(k+1)}(t;x,\Psi){\rm d}t\stackrel{{\scriptstyle\text{d}}}{{=}}A^{(k)}_{\lambda_{h}}\left[g(\cdot+S_{i});D_{i},\Psi\left(\cdot,\cdot+S_{i}\right)\right]$ (35)
@@
$A^{(k+1)}_{\lambda_{h}}(g;x,\Psi)-\int_{0}^{x}g(x-t)\mathrm{d}t$ $\geq\sum_{i=1}^{\eta}\int_{I_{i}(\lambda_{h})}g\circ W_{\lambda_{h}}^{(k+1)}(t;x,\Psi){\rm d}t$ (36)

### v2

@@ Abstract
where $W_{x}(\cdot)$ is the workload process in the queue and $g(\cdot)$ is any lower semi-continuous function. We also demonstrate how to utilise these bounds under the additional assumption that $\lambda(\cdot)$ is periodic in order to study the asymptotics of the tail distribution of the cycle length, deriving stability conditions, integrability of the steady-state distribution, etc.
@@ 1 Introduction
$G(\cdot)\equiv G(\Psi)(\cdot)\equiv\Psi(\cdot,\infty)\,,$
@@
$\ H(\cdot)\equiv H(\Psi)(\cdot)\equiv\Psi(\infty,\cdot)\equiv\lim_{s\uparrow\infty}\Psi(s,\cdot)\,.$
Note that in the context of Definition 1, for each $i\in\mathbb{N}$, $S_{i}$ (resp. $Y_{i}$) is the service (resp. patience) time of the $i$’th customer. Thus, since $(S_{i},Y_{i})$ is distributed according to $\Psi(\cdot)$ and $(S_{1},Y_{1}),(S_{2},Y_{2}),\ldots$ is an iid sequence, then $G(\cdot)$ (resp. $H(\cdot)$) is the service (resp. patience) distribution of the $\text{M}_{t}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queue. In particular, when $\Psi(\cdot)$ has a product-form (i.e., $\Psi(s,y)\equiv G(s)H(y)$ for any $(s,y)\in(0,\infty)\times(0,\infty]$), this model is known in the literature as the $\text{M}_{t}/\text{G}/1+\text{H}$ queue (see the work by Bodas, Mandjes and Ravner [11] for a multi-server version of this model). If, in addition, $\lambda(\cdot)$ is a constant function, then this queue will become M/G/$1$+H queue which have already been studied by several authors, e.g., Baccelli, Boyer and Hebuterne [4], Bae and Kim [5], Perry and Asmussen [49], Boxma, Perry, and Stadje [13] and Boxma, Perry, Stadje and Zacks, [12].
@@ 1.1 General description of the problem
However, as illustrated by the lower drawing of Figure 1, if the customers are allowed to have finite patience times, then (1) becomes false. This observation leads to the following question: Is there some sense (weaker than the one in (1)) in which $W_{\lambda}$ is stochastically dominated by $W_{\lambda_{h}}$? In the present work we provide an affirmative answer for this query by specifying a general class of functionals $\mathcal{H}\equiv\{h:\mathcal{D}([0,\infty))\rightarrow[0,\infty];h\in\mathcal{H}\}$ such that
@@
where $\mathcal{D}([0,\infty))$ is the Skorokhod space of all non-negative cádlág functions defined on $[0,\infty)$ and $\leq_{\text{st}}$ stands for a first-order stochastic dominance between the distributions of two random variables (see, e.g., the book by Shaked and Shanthikumar [45]).
@@ 1.2 The main results
$A_{\lambda}(g)\equiv A_{\lambda}(g;x)\equiv A_{\lambda}(g;x,\Psi)\equiv\int_{0}^{\tau_{\lambda}(x;\Psi)}g\circ W_{\lambda}(t;x,\Psi){\rm d}t\,,$
@@
i.e., the set of all non-negative lower semi-continuous functions on $[0,\infty)$. The following Theorems 1 and 2 are the main results of this work, as they jointly construct a general class of functionals for which (2) is valid.
@@
$A_{\lambda}(g;x,\Psi)\leq_{\text{st}}A_{\lambda_{h}}(g;x,\Psi),\ \forall x>0,\ g\in\mathcal{G},\ \Psi\in\mathcal{L}.$
@@
$A^{*}_{\lambda}(g;x,\Psi)\leq_{\text{st}}A^{*}_{\lambda_{h}}(g;x,\Psi),\ \forall x>0,\ g\in\mathcal{G},\ \Psi\in\mathcal{L}\,.$
@@
1.3 The Driving Force Behind the Proofs
This section includes a general description of the main benchmarks in the proofs of the main results. In addition, we provide some explanations of these benchmarks and discuss how they jointly yield the main results. The first step in the proof of Theorem 1 is to define the following auxiliary queue:
Definition 2
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    1. (a)
      • There is an empty place for him, i.e., the left limit of the queue-length process (counting also the potential customer in the service position) at his arrival epoch is less than $k+1$.
    2. (b)
Remark 1
$\tau^{(k)}_{\lambda}(x)\equiv\tau^{(k)}_{\lambda}\left(x;\Psi\right)\equiv\inf\left\{t\geq 0;W^{(k)}_{\lambda}\left(t;x,\Psi\right)=0\right\},$
and also for any Borel function $g:[0,\infty)\rightarrow[0,\infty)$
$A^{(k)}_{\lambda}(g)\equiv A^{(k)}_{\lambda}(g;x)\equiv A^{(k)}_{\lambda}(g;x,\Psi)\equiv\int_{0}^{\tau^{(k)}_{\lambda}(x;\Psi)}g\circ W^{(k)}_{\lambda}(t;x,\Psi){\rm d}t\,,$
The following lemma is the key to the proof of the main results.
Proposition 1
Assume that $\lambda(t)\leq\lambda_{h}$ for every $t\geq 0$. Then, for every Borel function $g:[0,\infty)\rightarrow[0,\infty)$, $x>0$, $\Psi\in\mathcal{L}$ and $k\in\mathbb{Z}_{\geq 0}$, the stochastic dominance $A^{(k)}_{\lambda}\left(g;x,\Psi\right)\leq_{\text{st}}A^{(k)}_{\lambda_{h}}\left(g;x,\Psi\right)$ holds and $A^{(k)}_{\lambda_{h}}\left(g;x,\Psi\right)$ is finite $\mathbb{P}$-a.s.
@@ 1.4 Periodic $\text{M}_{t}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queues and applications
$\xi_{\lambda}\equiv\inf\left\{\kappa n;\ W_{\lambda}(\kappa n)=0,\ n\in\mathbb{N}\right\}\,.$
@@
$\mathcal{A}^{*}_{\lambda}(g)\equiv\mathcal{A}^{*}_{\lambda}(g;\Psi)\equiv\int_{0}^{\xi_{\lambda}}g\circ W_{\lambda}(t-){\rm d}N_{\lambda}^{*}(t)\,.$
@@
Assume that $g\in\mathcal{G}$, $\Psi\in\mathcal{L}$ and $\lambda(\cdot)$ has a period $\kappa>0$ such that $\lambda(t)\leq\lambda_{h}\in(0,\infty)$ for every $t\geq 0$. In addition, let
$F_{\lambda_{h}}(\cdot)\equiv F_{\lambda_{h}}(\cdot;g,\Psi)\equiv{\text{dist}}\left[A_{\lambda_{h}}(g;x,\Psi)\right],$
$F^{*}_{\lambda_{h}}(\cdot)\equiv F^{*}_{\lambda_{h}}(\cdot;g,\Psi)\equiv{\text{dist}}\left[A_{\lambda_{h}}^{*}(g;x,\Psi)\right],$
when the initial distribution of $x$ is $G(\cdot)\equiv G(\Psi)(\cdot)$. In addition, denote the next two geometric decompoundings of $F_{\lambda_{h}}(\cdot)$ and $F^{*}_{\lambda_{h}}(\cdot)$:
$J_{\lambda_{h}}(u;g,\Psi)\equiv\sum_{n=1}^{\infty}\left(1-e^{-\kappa\lambda_{h}}\right)^{n-1}e^{-\kappa\lambda_{h}}F^{\star n}_{\lambda_{h}}(u-\max\{1,\kappa\}),\ \forall u\in\mathbb{R}\,,$
$J^{*}_{\lambda_{h}}(u;g,\Psi)\equiv\sum_{n=1}^{\infty}\left(1-e^{-\kappa\lambda_{h}}\right)^{n-1}e^{-\kappa\lambda_{h}}\left(F^{*}_{\lambda_{h}}\right)^{\star n}(u-\max\{1,\kappa\}),\ \forall u\in\mathbb{R}\,.$
@@
${\text{dist}}\left[\mathcal{A}_{\lambda}(g;\Psi)\right]\leq_{\text{st}}J_{\lambda_{h}}(\cdot;g,\Psi)\ \ \text{and}\ \ {\text{dist}}\left[\mathcal{A}^{*}_{\lambda}(g;\Psi)\right]\leq_{\text{st}}J^{*}_{\lambda_{h}}(\cdot;g,\Psi)\,.$
@@
$\lim_{u\to\infty}\sup\frac{\mathbb{P}\left\{\xi_{\lambda}>u\right\}}{1-G\left\{[u-g(0)\max\{\kappa,1\}](1-\rho_{h})\right\}}\leq\frac{1-e^{-\kappa\lambda_{h}}}{(1-\rho_{h})e^{-\kappa\lambda_{h}}}\,.$
@@
De Meyer and Teugels [15] proved that (8) is valid whenever $G(\cdot)$ is regularly-varying. Later, Zwart [57] extended this result to the class of all intermediate regularly-varying. It turns out that (8) is valid also when the tail of $G(\cdot)$ is heavier than $e^{-\sqrt{u}}$ (for the precise details see the works by Jelenković, and Momčilović [29] and Baltrūnas, Daley and Klüppelberg [6]). Another related work is by Asmussen, Klüppelberg and Sigman [3], who explained why (8) cannot hold when the tail of $G(\cdot)$ is lighter than $e^{-\sqrt{u}}$.
@@
$\lim_{u\to\infty}\sup\frac{\mathbb{P}\left\{\eta^{*}_{\lambda}>u\right\}}{1-G\left\{\frac{\left[u-g(0)\max\{\kappa,1\}\right](1-\rho_{h})}{\lambda_{h}}\right\}}\leq\frac{1-e^{-\kappa\lambda_{h}}}{(1-\rho_{h})e^{-\kappa\lambda_{h}}}\,.$
@@
Note that the fact that the $\text{M}_{t}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queue is stable under the assumption that $\Psi(\cdot)$ has a product-form and (10) is valid can also be proven by applying Lyapunov functions; see, e.g., Theorem 4.1 (and its proof) in the recent work by Bodas, Mandjes, and Ravner [11].
@@
$\mathbb{E}W_{\lambda}(\infty)=\frac{\mathbb{E}\mathcal{A}_{\lambda}(g)}{\mathbb{E}\xi_{\lambda}}\,,$
@@ 2 Proofs of Theorems 1 and 2
In this section we introduce the proofs of the main results of the present work (Theorem 1 and Theorem 2). Since the proofs of these two theorems are essentially identical (up to a few minor modifications), we will only provide a detailed proof of Theorem 1.
@@ 2.1 Proof of Proposition 1
As announced in Section 1.3, we use an inductive argument.
@@
$W_{\lambda}^{(0)}(t;x,\Psi)=W_{\lambda_{h}}^{(0)}(t;x,\Psi)=x-t\ \ ,\ \ \forall\ t\geq 0\,,x>0\ ,\ \Psi\in\mathcal{L}\,,$
@@
$\int_{I_{i}}g\circ W_{\lambda}^{(k+1)}(t;x,\Psi){\rm d}t\stackrel{{\scriptstyle{\text{d}}}}{{=}}A^{(k)}_{\lambda(T_{i}+\cdot)}\left[g(\cdot+S_{i});D_{i},\Psi\left(\cdot,\cdot+S_{i}\right)\right]$
for $1\leq i\leq\eta$. Note that (11) is based on the fact that the initial customer receives all his service requirement with probability one. To explain this fact, recall that for each $1\leq i\leq\eta$, $c_{i}$ initiates a busy period of an $\text{M}_{t}/\text{G}\big{(}\widetilde{\Psi}_{i}\big{)}/1/k+\text{H}\big{(}\widetilde{\Psi}_{i}\big{)}$ queue (for some $\widetilde{\Psi}_{i}$ to be specified later) that starts at his arrival epoch and lasts until he leaves the queue. Thus, the induction hypothesis (with $g\equiv 1$) implies that $I_{1},I_{2},\ldots,I_{\eta}<\infty$, $\mathbb{P}$-a.s. Therefore, since $\eta<\infty$, $\mathbb{P}$-a.s., the initial customer gets all his service requirement with probability one justifying the first term (i.e., the deterministic integral) in the RHS of (11).
Now, for each $1\leq i\leq\eta$ fix values $(d,s,t)\in\mathcal{S}\equiv\text{Supp}\left(D_{i},S_{i},T_{i}\right)$ and let $F(\cdot;d,s,t)$ (resp., $G(\cdot;d,s)$) be the cdf of
$A^{(k)}_{\lambda(t+\cdot)}\left[g(\cdot+s);d,\Psi\left(\cdot,\cdot+s\right)\right]\ \ \left(\text{resp., }A^{(k)}_{\lambda_{h}}\left[g(\cdot+s);d,\Psi\left(\cdot,\cdot+s\right)\right]\right).$
@@
$V_{i}(d,s,t)\equiv F^{-1}\left(U_{i};d,s,t\right),\qquad Z_{i}(d,s)\equiv G^{-1}\left(U_{i};d,s\right),$
@@
$V_{i}\equiv\begin{cases}V_{i}\left(D_{i},S_{i},T_{i}\right),&1\leq i\leq\eta,\\ 0,&\text{otherwise},\end{cases}\qquad\ Z_{i}\equiv\begin{cases}Z_{i}\left(D_{i},S_{i}\right),&1\leq i\leq\eta,\\ 0&\text{otherwise},\end{cases}$
@@
$\sum_{i=1}^{\eta}V_{i}\stackrel{{\scriptstyle\text{d}}}{{=}}\sum_{i=1}^{\eta}\int_{I_{i}}g\circ W_{\lambda}^{(k+1)}(t;x,\Psi){\rm d}t$
@@
$\sum_{i=1}^{\eta}\int_{I_{i}}g\circ W_{\lambda}^{(k+1)}(t;x,\Psi){\rm d}t\leq_{\text{st}}\sum_{i=1}^{\eta}Z_{i}\,,$
@@
$\int_{I_{i}(\lambda_{h})}g\circ W_{\lambda_{h}}^{(k+1)}(t;x,\Psi){\rm d}t\stackrel{{\scriptstyle\text{d}}}{{=}}A^{(k)}_{\lambda_{h}}\left[g(\cdot+S_{i});D_{i},\Psi\left(\cdot,\cdot+S_{i}\right)\right]$
@@
$A^{(k+1)}_{\lambda_{h}}(g;x,\Psi)-\int_{0}^{x}g(x-t)\mathrm{d}t$ $\geq\sum_{i=1}^{\eta}\int_{I_{i}(\lambda_{h})}g\circ W_{\lambda_{h}}^{(k+1)}(t;x,\Psi){\rm d}t$